Grade 12th passOrganic Chemistry

At NTP the viscosity co efficient of oxygen 0.2 milipoise.Find the collisiok diameter of oxygen.

Shreya
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Viscosity coefficient formulae is given as:
n = mv/(3root(2)*d^2)
v is the avg. velocity.
v = root(8RT/pi*M) => root(8*0.0821*273/3.14/32) = 1.335 m/s
on solving further and putting v in n we get,
d^2 = 50.35
d = 7.09 m.